OpenStudy (anonymous):

Any one wanna help with some alg 2 questions? fan and medal

OpenStudy (anonymous):

@dan815 @ganeshie8 @texaschic101 @rockstar0765 @paki

pooja195 (pooja195):

What 2 questions?

OpenStudy (anonymous):

What is the solution of log4x − 1032 = 5? and its Algebra 2 like the subject lol @pooja195

OpenStudy (owlcoffee):

You'll have to apply logarithmic properties n order to make that simpler.

OpenStudy (anonymous):

yeah i dont remember how to and i was wondering if someone would walk through it with me

OpenStudy (sweetburger):

alright is that the logbase 4 of x? or log base 10 of 4x

OpenStudy (anonymous):

i feel like its log base 4 cause thats exactly how it was written

OpenStudy (sweetburger):

sorry its just hard to tell on the computer :/

OpenStudy (anonymous):

its fine!

OpenStudy (sweetburger):

alright if it is log base 4 of x then we could say that 4^1037=x which seems very strange

OpenStudy (sweetburger):

that number is going to be massive

OpenStudy (anonymous):

yeah it seems massive already lol

OpenStudy (anonymous):

where did 1037 come from though?

OpenStudy (sweetburger):

1032+5 = 1037 i just moved the 1032 to the otherside

OpenStudy (anonymous):

oh okay i gotcha

OpenStudy (anonymous):

wouldnt it be -1032 +5?

OpenStudy (sweetburger):

alright so the equation says log4x -5=1032 so I added 5 to both sides so it became log4x = 1037. Also could u possible take a screenshot or pictures of the question because i feel like its not being expressed correctly the way it is written.

OpenStudy (anonymous):

nononono the equation is log4x-1032=5

OpenStudy (ybarrap):

It's more likely that the problem is $$ \Large { \ln^4x=5+1032=1037\\ \ln x=\sqrt[4]{1037}\\ x=\exp^{\sqrt[4]{1037}} } $$ Which is, not so massive. http://www.wolframalpha.com/input/?i=%281037%29%5E%281%2F4%29